### v1

@@ Discussion
For the benchmark data, training datasets were obtained by selecting all the clusters from three randomly chosen point clouds out of the 50 available. The remaining 47 were used for testing. A total of (non-blinking) or (blinking) point clouds were generated for the training of each scenario.
Metrics for performance evaluation
@@
Clustering in SMLM presents several challenges that impact the precise quantification of molecular organization and, consequently, the biological insights derived from these data. The effectiveness of a clustering algorithm should be evaluated based on its ability to accurately quantify key parameters, including the number of clusters, their positions, and the number of particles within each cluster.
@@
In our study, we report results obtained using several metrics, as summarized in Table 1, which extend beyond those used in the benchmark study [23]. In particular, we introduce cluster-level metrics, such as the Jaccard Index for cluster detection (JI), the root mean squared relative error in the number of localizations per cluster (RMSREN), and the root mean squared error in cluster centroid position (RMSEx,y). For their calculation, we first perform a distance-based pairing between clusters of the ground truth and prediction partitions using a Hungarian algorithm [45]. A predicted cluster is considered a true positive (TP) if its centroid is within a threshold distance from that of a ground truth cluster. It is considered a false positive (FP) if it has no corresponding cluster in the ground truth within . Similarly, a ground truth cluster with no corresponding predicted cluster within is accounted for as a false negative (FN).
We calculate the JI as
@@
Comparison of ARI-like metrics
@@
Simulated data
MIRO was validated on some of the datasets described in Ref. [23], with clustered localizations generated in 20002000 nm2 regions. Specifically, we used:
@@
C-shaped clusters: 64006400 nm2 images, with 30-60 clusters per image obtained by randomly placing localizations on semicircles with a radius of 250 nm and a radial standard deviation of 50 nm. Each cluster had a random number of localizations between 30 and 60, drawn from a uniform distribution. The number of non-clustered molecules was 6% of the total number of localizations, corresponding to 73% of the total number of structures (i.e., the sum of clusters and non-clustered localizations).
Ring-shaped clusters: 64006400 nm2 images, with 60-70 clusters per image obtained by randomly placing localizations on circles with a radius of 250 nm and a radial standard deviation of 50 nm. Each cluster had a random number of localizations between 60 and 80, drawn from a uniform distribution. The number of non-clustered localizations was 7% of the total number of localizations, corresponding to 84% of the total number of structures.
@@
For the training of MIRO’s model used for the analysis of integrin organization, we simulated groups of localizations distributed according to a 2-dimensional normal distribution with a standard deviation of 25-40 nm within 1000010000 nm2 regions. The number of localizations per cluster was drawn from a geometric distribution with an average of 25. The number of non-clustered localizations was 4% of the total number of localizations, corresponding to 60% of the total number of structures.
@@
For the proof of principle of the multiscale clustering, symmetric clusters having a random number of localizations between 6 and 30, drawn from a uniform distribution, with positions drawn from a 2-dimensional normal distribution with a standard deviation of 15 nm, were arranged with a 5 to 9-fold symmetry along a circle of radius 40 nm to form 40-60 rings. The number of non-clustered localizations was 1.5% of the total number of localizations, corresponding to 63% of the total number of structures. Images were 12501250 nm2.
@@
For the training of MIRO for the analysis of the NPC, we generated 12501250 nm2 images containing 5-9 NPC-like structures. Each structure was composed of 8 corners with a common vertex. For distributing the localizations, each corner was approximated as an isosceles triangle with height , which defined the circle radius as 50 nm. A random number of localizations from a uniform distribution between 0 and 80 were placed according to a normal distribution centered in the triangle centroid, with an angle-dependent standard deviation given by the center-to-perimeter distance divided by 1.8, providing an effective inner and outer radius of 20 and 80 nm, respectively. The number of non-clustered localizations was 3% of the total number of localizations, corresponding to 89% of the total number of structures.
Experimental data
@@
The experimental dataset used for the analysis of integrin organization was obtained by dSTORM imaging of a HeLa cell line modified to express ITGA5-HaloTag. In brief, cells were genetically edited with CRISPR/Cas9 to insert the HaloTag coding sequence at the 3’ terminus of ITGA5 cDNA. HeLa cells were co-transfected with a Cas9 and guide RNA-containing plasmid (lentiCRISPRv2 vector) and a repair plasmid (pBluescript) carrying the ITGA5-HaloTag cDNA, using polyethylenimine as the transfection reagent. Two days post-transfection, cells were treated with 1 g/ml puromycin for three days to select transfected cells, as lentiCRISPRv2 includes a puromycin resistance gene. Surviving cells were subsequently expanded for two weeks and subjected to three rounds of cell sorting.
@@
Cells () were seeded on fibronectin-coated (10 g/ml) glass-bottom dishes and allowed to adhere overnight. For staining, cells were incubated in a growth medium containing 200 nM Janelia Fluor® 647 HaloTag® Ligand for 30 minutes at 37 ∘ C, followed by three washes with fresh growth medium. An additional 1-hour incubation allowed for the removal of unbound ligand, followed by another three washes. Cells were then fixed in 2% paraformaldehyde for 20 minutes at room temperature.
@@
Single-molecule fluorescence imaging was carried out in a STORM imaging buffer with an inverted microscope (DMi8, Leica) equipped with a TIRF-illumination module (Infinity TIRF High Power, Leica) and a CMOS camera (Photometrics 95B). The beam of a 638-nm laser diode (180 mW, LBX-638, Oxxius) was combined with the beam of a 405-nm laser diode (50 mW, LBX-405, Oxxius) and coupled into the microscope through a polarization-maintaining monomode fiber. A 100 objective with a numerical aperture of 1.47 (HC PL APO 100/1.47, Leica) was used for TIRF illumination. The excitation beam was reflected into the objective by a quad-line dichroic beamsplitter and the fluorescence was detected through a quadruple band pass filter (set TRF89902 ET, Chroma). Photoactivation was manually controlled by the output power of the 405 nm laser and applied in adequate pulses. Fluorescence imaging was performed by excitation at 638 nm (0.2–0.5 kW/cm2). The camera was operated at a frame rate of 27 Hz.
@@
Images were processed with ThunderSTORM [48]. Data were fitted with a symmetric Gaussian PSF model using maximum likelihood estimation. The and localization positions were corrected for residual drift by an algorithm based on cross-correlation. Localizations were filtered by the localization precision ( 30 nm) and to exclude dim and very bright localizations ( counts ). Localizations persistent over consecutive frames detected within 40 nm from one another were merged into one localization.
Experimental data for the NPC were obtained from Ref. [13]
Acknowledgements
@@
JP acknowledges funding from the Adlerbertska Research Foundation (Grant AF2024-0305). GV acknowledges funding from the Horizon Europe ERC Consolidator Grant MAPEI (Grant No. 101001267), and the Knut and Alice Wallenberg Foundation (Grant No. 2019.0079). CM acknowledges support through grant RYC-2015-17896 funded by MCIN/AEI/10.13039/501100011033 and “ESF Investing in your future”, grants BFU2017-85693-R and PID2021-125386NB-I00 funded by MCIN/AEI/10.13039/501100011033/ and “ERDF A way of making Europe”, and grant AGAUR 2017SGR940 funded by the Generalitat de Catalunya.
Conflict of interest/Competing interests
@@
J.P., M.G., G.V., and C.M. hold shares and/or stock options of the company IFLAI AB. The other authors declare no competing interests.
Code availability
The MIRO software used in this article is available in the following repository: https://github.com/DeepTrackAI/MIRO/.

### v2

@@ 5.4 MIRO training and augmentations
For the benchmark scenarios, training datasets were constructed using all clusters from just three randomly selected point clouds out of the 50 available, with the remaining 47 used for testing. From the clusters contained in these point clouds, we typically generated (non-blinking) or (blinking) augmented training point clouds per scenario.
5.5 Ablation study and hyperparameter selection
To evaluate the robustness of MIRO and guide its configuration, we performed a series of ablation studies targeting its core hyperparameters and architectural components. These analyses aim to determine the sensitivity of clustering performance to specific design choices and to identify suitable parameter settings across diverse datasets. A summary of the hyperparameters used for each scenario is provided in Supplementary Table 7.
5.5.1 Number of recurrent steps: influence on performance and oversmoothing
MIRO adopts a recurrent architecture in which a single graph transformation block is applied multiple times. As previously emphasized, the same set of weights is reused at each step, so increasing the number of recurrent steps does not affect the number of trainable parameters. Instead, this depth controls the effective receptive field of each node, determining how far information can propagate across the graph.
Even in scenarios with relatively simple structure (e.g., a single clustering level), a sufficient number of recurrent steps may be required to enable long-range interactions.
To demonstrate this effect, we conducted an ablation study on the Scenario 6 with blinking, varying the number of recurrent steps from 5 to 25. We quantified clustering quality using the compression index, defined as:
where, denotes the number of clusters, the set of indices for cluster , the position of point , its predicted displacement, and the centroid of cluster . A value of compression index close to 1 indicates that the predicted displacements successfully bring the points close to their respective cluster centers (i.e., high compression), whereas a value near 0 reflects little to no compression. This metric quantifies the network’s ability to contract clustered localizations toward a common center.
As shown in Supplementary Figure 2a, the compression index calculated for MIRO consistently increases with increasing depth, flattening around 15 recurrent steps. These results (mean ± standard deviation over 5 runs) indicate that deeper recurrent iterations improve feature integration and enhance clustering compactness.
This behavior also highlights a critical difference between MIRO and conventional message-passing architectures. Classical message-passing GNNs are prone to oversmoothing [43], a phenomenon in which node features become indistinguishable as the number of layers increases. This issue is commonly diagnosed using the Dirichlet energy:
where denotes the feature vector of node at the final recurrent step , and is the total number of nodes in the graph [50].
Supplementary Figure 2a–b compares MIRO to standard message-passing networks of similar depth on the same dataset. While MIRO achieves consistent compression index values and maintains high Dirichlet energy beyond 15 recurrent steps, traditional GNNs exhibit rapid oversmoothing, with both metrics deteriorating as the network deepens. This underscores MIRO’s ability to preserve feature diversity and spatial coherence, even with deep processing.
5.5.2 Dimensionality of the hidden representation
To assess the appropriate model capacity for MIRO, we performed an ablation study on the dimensionality of the hidden node and edge features.
Our objective was to identify the smallest hidden dimension that consistently guarantees high clustering performance across all tested scenarios. To this end, we selected a challenging synthetic dataset containing mixed shapes (spots and ellipses), where the network must simultaneously learn localization patterns and distinguish between structural classes.
We evaluated dimensionalities of 64, 128, 256, and 512 to identify an optimal hidden size for MIRO’s node and edge representations. As shown in Supplementary Table 8, a dimensionality of 256 yields the best overall performance across a broad set of metrics.
Smaller hidden dimensions, such as 64 and 128, result in a minor drop in both clustering quality and spatial resolution. However, performance remains respectable even at 64 dimensions, as the essential structural features are preserved and metric scores remain competitive.
Conversely, increasing the dimensionality to 512 does not lead to significant improvements. In fact, the gains plateau or slightly regress for most metrics, while computational cost and memory usage increase.
These results support the use of a 256-dimensional hidden representation as a robust default configuration for diverse tasks, offering a favorable balance between expressive power and efficiency. They also reinforce the idea that MIRO achieves its performance not through brute-force parameter scaling, but through its effective recurrent architecture and well-structured loss function.
5.5.3 Loss function weights
MIRO’s loss function consists of multiple components, depending on the task. In all cases, the core objective includes two terms: a displacement loss , which encourages nodes to move toward their cluster centers, and a distance-preserving auxiliary loss , which maintains relative distances between neighboring points after displacement. The contribution of the latter is modulated by the hyperparameter (Eq. 4).
In classification scenarios, such as those involving multiple structural types, a third term is added: a categorical cross-entropy loss , weighted by the hyperparameter , which guides the model in predicting shape labels for individual nodes.
To evaluate the sensitivity of performance to these hyperparameters, we conducted ablation studies on a representative dataset. Specifically, we selected a multishape dataset combining circular and elliptical clusters, as it provides a comprehensive test case with diverse structural features.
The results in Supplementary Table 9 demonstrate that MIRO maintains consistently high performance across a broad range of hyperparameter values. For the clustering loss weight , values between 0 and 20 yield comparable results in terms of overall clustering quality, with intermediate settings providing marginal gains in compactness and separation metrics such as ARI and IoU.
At first glance, this robustness might suggest that the auxiliary loss weighted by is unnecessary. However, its role becomes critical in challenging conditions. As shown in Supplementary Figure 3, when applied to datasets affected by blinking artifacts, setting is essential to achieve a stable compression index. This highlights the utility of the auxiliary term in improving robustness under noisy conditions, while also confirming that MIRO’s core architecture remains effective even in its absence.
Regarding the classification loss weight , performance remains stable for values between 0.05 and 0.5, with F1 scores consistently above 0.92, indicating reliable structural classification. Moderate values of (e.g., 0.15–0.2) yield slight improvements in ARI and RMSEx,y, suggesting that a moderate emphasis on classification enhances spatial precision without compromising clustering performance.
Overall, these findings confirm that MIRO is not highly sensitive to the exact choice of and , simplifying its configuration and highlighting its robustness across multiple simultaneous tasks.
5.5.4 Training efficiency and practical considerations
Training time is not substantially affected by the number of recurrent steps, as the model’s capacity remains constant, with a total of 528,386 trainable parameters. On an NVIDIA A100 GPU with 40 GB of available memory, training a MIRO model takes approximately 20 to 60 minutes per dataset, with variation primarily driven by dataset size and graph density. Memory usage remains modest, owing to the model’s shallow architecture and relatively small parameter count.
5.6 Metrics for performance evaluation
@@
Clustering in SMLM presents several challenges that impact the precise quantification of molecular organization and, consequently, the biological insights derived from these data. The effectiveness of a clustering algorithm should be evaluated based on its ability to accurately quantify key parameters, including the number of clusters, their positions, and the number of objects within each cluster.
@@
In our study, we report results obtained using several metrics, as summarized in Table 1, which extend beyond those used in the benchmark study [23]. In particular, we introduce cluster-level metrics, such as the Jaccard Index for cluster detection (JIc), the root mean squared relative error in the number of localizations per cluster (RMSREN), and the root mean squared error in cluster centroid position (RMSEx,y). For their calculation, we first perform a distance-based pairing between clusters of the ground truth and prediction partitions using a Hungarian algorithm [51]. A predicted cluster is considered a true positive (TP) if its centroid is within a threshold distance from that of a ground truth cluster. It is considered a false positive (FP) if it has no corresponding cluster in the ground truth within . Similarly, a ground truth cluster with no corresponding predicted cluster within is accounted for as a false negative (FN).
We calculate the JIc as
@@
5.6.1 Comparison of ARI-like metrics
@@
5.7 DBSCAN parameter selection
In the benchmark study [23], distinct DBSCAN parameter sets ( and minPts) were used to maximize ARI or IoU, independently. While this approach may be appropriate for assessing theoretical performance limits, it raises important practical concerns. Most notably, it becomes unclear which set of parameters should be adopted in real-world applications, as the optimal configuration for ARI may differ significantly from that for IoU — an issue clearly illustrated in the scenarios with blinking, where the two metrics are optimized by substantially different parameter choices (see Supplementary Table 1 in Ref. [23]).
Furthermore, although the benchmark reports only a single parameter pair per scenario and metric, the best metric values reported in their comparison appear to have been derived from an optimization procedure conducted over each FOV. Specifically, the authors refer to results in Figure 3 and 4 as the “average peak score” or “mean of the maximal ARI and IoU scores”, suggesting that the best metric values were obtained by tuning parameters independently for each FOV before averaging the results. This interpretation is supported by our own attempts to reproduce the reported values, which provided comparable results only when optimizing parameters at the level of individual FOVs, yielding a range of and minPts values across each scenario.
While such FOV-specific optimization may be informative for gauging the best-case performance of a method, it is not feasible in real experimental settings, where ground truth is unavailable. Moreover, it provides little guidance to practitioners on how to select suitable parameters for a new dataset.
To ensure a realistic and fair comparison, we adopted a consistent strategy: for each scenario, a single pair of DBSCAN parameters was applied uniformly across all FOVs. The selected parameters were those that maximized the ARI metric. For the benchmark datasets analyzed with DBSCAN, we used the DBSCAN settings that achieved the highest ARI scores as reported in Supplementary Table 1 in Ref. [23].
5.8 Simulated data
MIRO was validated on all the datasets (both without and with blinking) described in Ref. [23], with clustered localizations generated in 20002000 nm2 regions. For the results described in Figure 2 and Table 1, we used:
@@ 5.8 Simulated data
In Supplementary Table 3 and Supplementary Table 4, we report the results obtained for all scenarios, without and with blinking, respectively. The additional scenarios are:
Scenario 2: 20 clusters of 15 molecules per cluster with 50% of the total molecules being clustered.
Scenario 3: 20 clusters of 15 molecules per cluster with 20% of the total molecules being clustered.
Scenario 4: 20 clusters of 5 molecules per cluster with 25% of the total molecules being clustered.
Scenario 7: 10 clusters with a width of 25 nm and 10 clusters with a width of 75 nm, with 50% of the total molecules clustered.
@@
Scenario 10: 20 clusters of 15 molecules per cluster with 50% of the total molecules being clustered, but with non-clustered molecule density increasing from left to right across the region.
@@
C-shaped clusters: 64006400 nm2 images, with 30–60 clusters per image obtained by randomly placing localizations on semicircles with a radius of 250 nm and a radial standard deviation of 50 nm. Each cluster had a random number of localizations between 30 and 60, drawn from a uniform distribution. The number of non-clustered molecules was 6% of the total number of localizations, corresponding to 73% of the total number of structures (i.e., the sum of clusters and non-clustered localizations).
Ring-shaped clusters: 64006400 nm2 images, with 60–70 clusters per image obtained by randomly placing localizations on circles with a radius of 250 nm and a radial standard deviation of 50 nm. Each cluster had a random number of localizations between 60 and 80, drawn from a uniform distribution. The number of non-clustered localizations was 7% of the total number of localizations, corresponding to 84% of the total number of structures.
@@
For the training of MIRO’s model used for the analysis of integrin organization, we simulated groups of localizations distributed according to a 2-dimensional normal distribution with a standard deviation of 25–40 nm within 1000010000 nm2 regions. The number of localizations per cluster was drawn from a geometric distribution with an average of 25. The number of non-clustered localizations was 4% of the total number of localizations, corresponding to 60% of the total number of structures.
@@
For the proof of principle of the multiscale clustering, symmetric clusters having a random number of localizations between 6 and 30, drawn from a uniform distribution, with positions drawn from a 2-dimensional normal distribution with a standard deviation of 15 nm, were arranged with a 5 to 9-fold symmetry along a circle of radius 40 nm to form 40–60 rings. The number of non-clustered localizations was 1.5% of the total number of localizations, corresponding to 63% of the total number of structures. Images were 12501250 nm2.
@@
For the training of MIRO for the analysis of the NPC, we generated 12501250 nm2 images containing 5–9 NPC-like structures. Each structure was composed of 8 corners with a common vertex. For distributing the localizations, each corner was approximated as an isosceles triangle with height , which defined the circle radius as 50 nm. A random number of localizations from a uniform distribution between 0 and 80 were placed according to a normal distribution centered in the triangle centroid, with an angle-dependent standard deviation given by the center-to-perimeter distance divided by 1.8, providing an effective inner and outer radius of 20 and 80 nm, respectively. The number of non-clustered localizations was 3% of the total number of localizations, corresponding to 89% of the total number of structures.
5.9 Experimental data
@@
The experimental dataset used for the analysis of integrin organization was obtained by dSTORM imaging of a HeLa cell line modified to express ITGA5-HaloTag. In brief, cells were genetically edited with CRISPR/Cas9 to insert the HaloTag coding sequence at the 3’ terminus of ITGA5 cDNA. HeLa cells were co-transfected with a Cas9 and guide RNA-containing plasmid (lentiCRISPRv2 vector) and a repair plasmid (pBluescript) carrying the ITGA5-HaloTag cDNA, using polyethylenimine as the transfection reagent. Two days post-transfection, cells were treated with puromycin for three days to select transfected cells, as lentiCRISPRv2 includes a puromycin resistance gene. Surviving cells were subsequently expanded for two weeks and subjected to three rounds of cell sorting.
@@
Cells () were seeded on fibronectin-coated () glass-bottom dishes and allowed to adhere overnight. For staining, cells were incubated in a growth medium containing 200 nM Janelia Fluor® 647 HaloTag® Ligand for 30 minutes at 37 ∘C, followed by three washes with fresh growth medium. An additional 1-hour incubation allowed for the removal of unbound ligand, followed by another three washes. Cells were then fixed in 2% paraformaldehyde for 20 minutes at room temperature.
@@
Single-molecule fluorescence imaging was carried out in a STORM imaging buffer with an inverted microscope (DMi8, Leica) equipped with a TIRF-illumination module (Infinity TIRF High Power, Leica) and a CMOS camera (Photometrics 95B). The beam of a 638-nm laser diode (180 mW, LBX-638, Oxxius) was combined with the beam of a 405-nm laser diode (50 mW, LBX-405, Oxxius) and coupled into the microscope through a polarization-maintaining monomode fiber. A 100 objective with a numerical aperture of 1.47 (HC PL APO 100/1.47, Leica) was used for TIRF illumination. The excitation beam was reflected into the objective by a quad-line dichroic beamsplitter and the fluorescence was detected through a quadruple band pass filter (set TRF89902 ET, Chroma). Photoactivation was manually controlled by the output power of the 405 nm laser and applied in adequate pulses. Fluorescence imaging was performed by excitation at 638 nm (0.2–0.5 kW/cm2). The camera was operated at a frame rate of 27 Hz. Experiments were performed in three independent biological replicates.
@@
Images were processed with ThunderSTORM [54]. Data were fitted with a symmetric Gaussian PSF model using maximum likelihood estimation. The and localization positions were corrected for residual drift by an algorithm based on cross-correlation. Localizations were filtered by the localization precision ( nm) and to exclude dim and very bright localizations (counts). Localizations persistent over consecutive frames detected within 40 nm from one another were merged into one localization.
Experimental data for the NPC were obtained from Ref. [13].
5.10 Comparison with a supervised graph-based clustering framework
To compare MIRO with recent supervised graph-based clustering methods, we implemented a two-stage GNN pipeline inspired by the approach described in Ref. [22], hereafter referred to as MAGIK-S. Clustering performance was evaluated on two representative benchmark configurations (Scenarios 5 and 6), as reported in Supplementary Table 6.
MAGIK-S consists of two stages. In the first stage, a GNN classifies each localization as clustered or non-clustered, serving as a background filter. In the second stage, a separate GNN processes the filtered point cloud and predicts binary edge labels indicating whether each edge connects localizations belonging to the same cluster. These predictions are used to prune the graph, and final clusters are extracted using the Louvain community detection algorithm. Both GNNs employ the MAGIK architecture [31].
The input graph construction in Stage 1 matches that of MIRO, with the exception that absolute spatial coordinates are used as node features. Stage 2 operates only on localizations predicted as clustered in Stage 1.
To ensure a fair comparison, MAGIK-S was configured to use the same latent dimension (256) as MIRO and three Fingerprinting Graph Neural Network (FGNN) layers, each incorporating four attention heads. The training procedure was the one described in Ref. [22] whereas the dataset configuration followed the one used for MIRO.
Supplementary information Supplementary Information is available for this paper, containing three figure and eight tables.
@@ 5.10 Comparison with a supervised graph-based clustering framework
Acknowledgements J.P. acknowledges funding from the Adlerbertska Research Foundation (Grant AF2024-0305). G.V. acknowledges support from the Horizon Europe ERC Consolidator Grant MAPEI (grant number 101001267) and the Knut and Alice Wallenberg Foundation (grant number 2019.0079). C.M. acknowledges support through grant RYC-2015-17896 funded by MCIN/AEI/10.13039/501100011033 and “ESF Investing in your future”, grants BFU2017-85693-R and PID2021-125386NB-I00 funded by MCIN/AEI/10.13039/501100011033/ and “ERDF A way of making Europe”.
@@
Conflict of interest/Competing interests J.P., M.G., G.V., and C.M. hold shares and/or stock options of the company IFLAI AB. The other authors declare no competing interests.
Data availability The benchmark simulations from Ref. [23] used in this study are publicly available at https://github.com/DJ-Nieves/ARI-and-IoU-cluster-analysis-evaluation. The NPC dataset from Ref. [13] is publicly available at https://www.ebi.ac.uk/biostudies/BioImages/studies/S-BIAD8.
Code availability The code used to develop the model, perform the analyses, and generate the results of this study is publicly available and has been deposited in GitHub at https://github.com/DeepTrackAI/MIRO/, under MIT license. The specific version of the code associated with this publication is archived in Zenodo and is accessible at https://doi.org/10.5281/zenodo.17194652 [55].